OpenStudy (samsan9):

What are the coordinates of the focus of the conic section shown below? Y^2+16y+4x+4=0

OpenStudy (samsan9):

please help i have to know how to do this by monday

OpenStudy (anonymous):

Do you know how to complete the square?

OpenStudy (samsan9):

isnt that when it factors perfectly?

OpenStudy (anonymous):

yes, that's the idea. first you need to move the 4x + 4 to the right of the equation y^2 + 16y = -4x -4

OpenStudy (anonymous):

then complete the square y^2 +16y + 64 = -4x -4 + 64 (y+8)^2 = -4x +60 (y+8)^2 = 4 (-1) (x-15) Does this look familiar now?

OpenStudy (samsan9):

yes it sort of looks familiar

OpenStudy (anonymous):

Standard form is (y-k)^2 = 4p (x-h) Do you know which conics this is?

OpenStudy (anonymous):

Parabola opens sideways

OpenStudy (samsan9):

its in parabola form isn't it?

OpenStudy (anonymous):

Correct !

OpenStudy (samsan9):

ok so what is next?

OpenStudy (anonymous):

Use the textbook formula to find focus?

OpenStudy (anonymous):

which is( h+p, k)

OpenStudy (samsan9):

(-15+?,8),

OpenStudy (anonymous):

remember it's x- h and y - k in this case h is postivie and y is negative because the original equation is (x+8)^2 = (4) (-1) (x-15) ) so (15 + ?, -8) p is -1 (15-1,- 8)

OpenStudy (samsan9):

(14,-8) is the correct answer :o

OpenStudy (anonymous):

That's what I got too :D

OpenStudy (samsan9):

thank you for being awesome do you think you could help me with a few tomorrow?

OpenStudy (anonymous):

Sure, what time though? I have to go to church around 10 am to 4 pm tomorrow :P

OpenStudy (anonymous):

I live in california so pacific time

OpenStudy (samsan9):

ima try to wake up around 7 am pacific time

OpenStudy (anonymous):

That's fine! I usually wake up around 6 am so I'll be there :D

OpenStudy (samsan9):

ok ill fan you and thank you so much:D

OpenStudy (anonymous):

welcome! It's 11:42 pm here so good night, and see you tomorrow.

OpenStudy (samsan9):

alright you too